# APoW: Auditable Proof-of-Work Against Block Withholding Attacks

### v1

@@ Abstract
We introduce APoW , a novel proof-of-work (PoW) construction inspired by Hashcash-style nonce searching, which enables the auditing of other miners’ work through accountable re-scanning of the nonce space. The proposed scheme allows a miner to probabilistically attest to having searched specified regions of the nonce space in earlier mining rounds, while concurrently earning rewards for performing productive work for a new block or pool share. This capability enables miners belonging to a mining pools to audit another miner’s claimed effort retroactively, thereby allowing the probabilistic detection of block withholding attacks (BWAs) without requiring trusted hardware or trusted third parties. As a consequence, the construction supports the design of decentralized mining pools in which work attribution is verifiable and withholding incentives are substantially reduced. The scheme preserves the fundamental properties of conventional PoW, including public verifiability and difficulty adjustment, while adding an orthogonal auditability layer tailored to pool-based mining. Finally, while a full deployment of APoW in Bitcoin would require a consensus rule change and minor modifications to mining ASICs, the construction remains practically useful even without consensus changes, for instance, as a pool-level auditing mechanism that enables verifiable pay-for-auditing using existing pool reserves.
@@ 1 Introduction
BWAs are particularly difficult to detect because, under standard proof-of-work constructions, the absence of a block is not itself evidence of misbehavior. Block discovery is a low-probability event, and variance alone can explain extended periods without successful blocks even for honest miners. Consequently, distinguishing deliberate withholding from statistical fluctuation requires long observation windows and provides, in the best case, weak confidence guaranties. As a result, most existing pool protocols cannot attribute blame or provide cryptographic evidence of withholding behavior.
@@ 2.1 Rosenfeld’s oblivious shares proposal
While conceptually appealing, oblivious shares introduces several important constraints that are in tension with modern pooled-mining practice:
@@
    Hard fork / header reinterpretation. To apply an oblivious-shares-style mechanism to Bitcoin, a hard-fork is required: a difficult to achieve network-wide coordination procedure. Both full nodes and SPV/light clients would need to upgrade to preserve their security model, since the proof-of-work predicate they implicitly rely on is tied to header semantics. Alternative variants that avoid requiring light-client upgrades have been discussed, but at a cost of significantly reduced security for those clients [7, 6].
@@ 2.2 Rosenfeld’s “pop quiz” proposal
2.3 Peer Reviews and Short Lived Audits Proposal
@@
3 APoW Compared to Prior Work
We compare our scheme with the existent solutions to BWA.
3.1 Oblivious Shares vs Our Approach
3.2 “Pop quiz” (and its variants) vs Our Approach
In pop-quiz schemes, the hashing spend verifying other miners is not rewarded by the blockchain, limiting their practical effectiveness as a BWA deterrent.
@@ 4 Description
In Section 5, we describe how the search space $S_{i}^{A}$ associated with a miner $A$ is communicated to the mining pool and subsequently assigned to an auditor in a centralized pool setting. We further discuss how analogous information can be disseminated in a decentralized mining pool, allowing multiple participants to independently reason about and audit the claimed search space $S_{i}^{A}$.
@@ 4.1 Scheme 1: Non-Empty V-Blocks
Let us assume that we associate a new block header with the PoW after it is found. Then the parent block will also be chosen post-PoW, which means that a v-block could be sealed to fit in different slots of the blockchain. This could disrupt the incentives of mining. A careful analysis is required including how this impact mining fairness and selfish mining attacks. Without signature equivocation and penalization of a stake, generating valid v-blocks is free. To prevent attacks, we add a blockchain consensus rule that a v-block can be included only just immediately after its audited block. Without this consensus rule, the protocol would need to check each new v-block against all past v-blocks to prevent double-inclusion. The downside of the immediate inclusion rule is that the audit time is reduced to a single block interval, which can vary widely.
@@ 4.2 Scheme 2: Empty V-Blocks
When a v-block does not carry neither transaction nor block timestamp information, the v-block does not need sealing and is already unique. Although this option is possible, it has the downside that it reduces the blockchain transaction throwout. We analyze this variant in Appendix B.
@@ 4.3 Fairness
Although our scheme is most effective when the underlying blockchain protocol explicitly rewards the submission of verification blocks (v-blocks), such rewards are not strictly required for the construction to remain useful. Even in the absence of direct compensation for hash power devoted to auditing, a mining pool can still safely assign resources to verification tasks: shares generated during auditing are indistinguishable from ordinary shares and can therefore be accounted for under the pool’s existing payout mechanisms. As long as the expected cost of auditing remains lower than the expected loss induced by a successful block withholding attack, the pool can rationally fund audits without exposing itself to additional avenues of cheating.
@@ 4.7 Avoidance of Long Forks
Figure 2: A blockchain mixing blocks and v-blocks auditing previous work units (w.u.)
@@ 4.8 V-Block Attribution
Every block $i$ contains the field $\mathsf{poolAddress}_{i}$ in its block template. This is the address where v-blocks are paid. The address is extracted from the message $M$ in the v-block. The address can be controlled by the pool server, or it can be a smart contract controlled address in the mainchain (if it has this capability) or it can be a sidechain exit address that forwards the funds to a smart contract on a sidechain. The pool sever or pool smart contract is responsible for sharing the reward among the workers, in the same manner as a normal block reward. The server can also track who was the auditor of the audited block and pay a special prize to that worker. Alternatively, a field $\mathsf{lastBlock}_{i}$ could be added to the block template. This field would indicate the block ID of the last block created by the same pool that was included in the canonical chain. The protocol could retrieve that block coinbase transaction and distribute the reward to the same addresses and in the same proportion used in the referenced block. This enables Bitcoin decentralized mining pools that pay their workers directly from the coinbase transaction using multiple outputs.
@@
4.9 Pool Shares
@@
4.10 Attacks
There are two potential attacks related to normal block mining and two related to v-mining.
@@ 4.10 Attacks
Memory-CPU trade-off provides little help with regard to RAM costs. Assuming that all hash operations are stored sequentially, each cached hash operation $j$ requires the storage of $X_{j}$, which is 256-bits in length. It is neither necessary nor optimal to store $X_{j}$ in full. The digests $X_{j}$ could be partially stored and recomputed in full after a positive match with the partial target pattern, to verify the full match. For example, if we store 16-bits per digest, we can retrieve 8 hash results per RAM memory access (assuming a 128-bit wide data bus). This results in a 8X speedup for RAM access and still leaves plenty of time for a CPU or a non-pipelined hashing core to recompute promising partial matches. Since the nonce candidates to check will not be sequential, they need to be transmitted to the hashing component, thus the communication can become the system’s bottleneck. While this optimization could make storing and retrieving results as fast as hashing, the v-mining machine would still cost over 1T dollars.
The attacker can also store just 1 bit per hash, packing bits in a 128-bit hint word that can be retrieved in a single RAM access. The attack requires the building of a special ASIC architecture that reads hint words, scans its bits while it increments the nonce counter at double clocking speeds. This small overclocked circuit would stall for 1 cycle when a valid bit (0 or 1) is found in hint bit sequence, so that the corresponding nonce enters the hashing pipeline, keeping it almost full. Although this design may work, the machine would still cost 39B USD due to its large RAM requirement.
@@
Figure 3: Recursive Auditing: audit at height $j$ conducted for a auditing work unit at height $k$, which in turns audits a work unit at height $i$
@@
4.11 Compatibility with Oblivious Shares
@@ 4.13 Migrating Bitcoin PoW to APoW
While switching to a new mining function and disposing all pre-existing ASICs is unfeasible, Bitcoin could hard-fork to enable both methods of mining to co-exists. Mining pools could offer an additional incentive for miners that allow v-mining, and with enough hashrate v-mining, no miner would require to pass KYC checks or amass reputation.
@@
5 Defining Auditable Work Units in Mining Pools
@@ Appendix B Unsealed V-Blocks
    Scheme 2c. The v-block header is not part of the blockchain header chain, but is included in the blockchain inside a normal transaction $T$ with a special payload (i.e., OP_RETURN data in Bitcoin, or calldata in EVM-like chains). The v-block does not contribute to the cumulative work of the chain, but the protocol mints a subsidy $S$ to reward the v-miner. The transaction $T$ has an output with value $S$, which is considered as an virtual input of the transaction, same as the coinbase.
@@
Mining v-blocks should be almost as profitable as mining normals blocks to avoid penalizing pools that use them to perform audits. However, it should not be equally profitable. If so, the mining pools could engage only in v-mining, reducing the available transaction volume on the blockchain and preventing other pools from auditing. In Bitcoin, v-mining would be discouraged as unsealed v-blocks cannot earn transaction fees.

### v2

@@ Abstract
We introduce Auditable Proof-of-Work (APoW), a novel proof-of-work (PoW) construction inspired by Hashcash-style nonce searching, which enables the auditing of other miners’ work through accountable re-scanning of the nonce space. The proposed scheme allows a miner to probabilistically attest to having searched specified regions of the nonce space in earlier mining rounds, while concurrently earning rewards for performing productive work for a new block or pool share. This capability enables miners belonging to a mining pools to audit another miner’s claimed effort retroactively, thereby allowing the probabilistic detection of block withholding attacks (BWAs) without requiring trusted hardware or trusted third parties. As a consequence, the construction supports the design of decentralized mining pools in which work attribution is verifiable and withholding incentives are substantially reduced. The scheme preserves the fundamental properties of conventional PoW, including public verifiability and difficulty adjustment, while adding an orthogonal auditability layer tailored to pool-based mining. Finally, while a full deployment of APoW in Bitcoin would require a consensus rule change and minor modifications to mining Application-Specific Integrated Circuits (ASICs), the construction remains practically useful even without consensus changes, for instance, as a pool-level auditing mechanism that enables verifiable pay-for-auditing using existing pool reserves.
Keywords: Proof-of-Work, Bitcoin, Auditable Proof-of-Work; APoW; Block Withholding Attacks; Mining Pools; Decentralized Mining; Hashcash; Nonce Searching
@@ 1 Introduction
BWAs are particularly difficult to detect because, under standard proof-of-work constructions, the absence of a block is not itself evidence of misbehavior. Block discovery is a low-probability event, and variance alone can explain extended periods without successful blocks even for honest miners. Consequently, distinguishing deliberate withholding from statistical fluctuation requires long observation windows and provides, in the best case, weak confidence guarantees. As a result, most existing pool protocols cannot attribute blame or provide cryptographic evidence of withholding behavior.
@@
1.1 Mining Pool Payment Schemes and Exposure to Block-Withholding Attacks
Mining pools employ a variety of reward distribution schemes to compensate participating miners for contributed hash power. These schemes differ significantly in how mining variance is allocated between miners and pool operators, and consequently in their susceptibility to block-withholding attacks (BWAs), in which a miner submits valid low-difficulty shares while deliberately withholding full-difficulty blocks.
Block-withholding incentives.
A key observation from prior work is that the profitability of a BWA depends critically on whether a miner’s payout is conditioned on the pool’s actual block discovery events. If rewards are independent of block discovery, a withholding miner can preserve their expected income while reducing the pool’s revenue, thereby creating a strictly profitable deviation.
Classification of payment schemes.
We classify the most widely deployed mining pool payment schemes according to their resistance to BWAs:
  • Block-conditioned schemes, such as Pay-Per-Last-$N$-Shares (PPLNS) and score-based variants, distribute rewards only when the pool successfully mines a block. In these schemes, withholding a valid block reduces the attacker’s own expected revenue, rendering BWAs economically irrational.
  • Share-conditioned schemes, such as Pay-Per-Share (PPS), compensate miners solely based on submitted shares, regardless of block discovery outcomes. These schemes transfer all variance risk to the pool operator and make BWAs strictly profitable.
  • Hybrid schemes, such as PPS+, partially condition rewards on block discovery (typically for transaction fees), resulting in intermediate exposure to BWAs.
  • Full Pay-Per-Share (FPPS) schemes extend PPS to both the block subsidy and transaction fees by paying miners the expected value of each per share, thereby maximizing pool-side variance and BWA exposure.
Security comparison.
Table 1 summarizes this classification.
Scheme Variance Bearer BWA Resistance BWA Profitability
PPLNS Miner Strong No
Score-based Miner Strong No
PPS+ Shared Moderate Partial
PPS Pool Weak Yes
FPPS Pool Very weak Yes
Table 1: Classification of mining pool payment schemes by resistance to block-withholding attacks.
Empirical exposure of Bitcoin hashrate.
Based on public disclosures by major mining pools, historical hashrate distribution data, and advertised payout options, we estimate that approximately 25%–35% of the current Bitcoin network hashrate participates in pools offering PPS or FPPS-style payouts, and is therefore economically susceptible to block-withholding attacks in the absence of additional detection or auditing mechanisms. This estimate accounts for large pools that either operate exclusively under PPS/FPPS or offer these schemes as a selectable option, and excludes pools that rely solely on PPLNS or score-based methods.
While the exact fraction varies over time with market conditions and miner preferences, the prevalence of PPS-style payouts highlights a fundamental tension between payout predictability and incentive compatibility in mining pool design.
Implications.
The above classification illustrates that widely deployed pool payment mechanisms remain vulnerable to BWAs unless miner compensation is explicitly tied to verifiable block discovery. This motivates the exploration of alternative constructions, such as auditable proof-of-work mechanisms, that can preserve predictable payouts while restoring incentive compatibility without requiring trusted pool operators or trusted hardware.
@@ 2.1 Rosenfeld’s Oblivious Shares Proposal
While conceptually appealing, oblivious shares introduce several important constraints that are in tension with modern pooled-mining practice:
@@
    Hard fork / header reinterpretation. To apply an oblivious-shares-style mechanism to Bitcoin, a hard-fork is required: a difficult-to-achieve network-wide coordination procedure. Both full nodes and SPV/light clients would need to upgrade to preserve their security model, since the proof-of-work predicate they implicitly rely on is tied to header semantics. Alternative variants that avoid requiring light-client upgrades have been discussed, but at a cost of significantly reduced security for those clients [7, 6].
@@ 2.2 Rosenfeld’s “pop quiz” Proposal
2.3 Peer Reviews and Short-Lived Audits Proposal
@@ 3 Description
In Section 6.5, we describe how the search space $S_{i}^{A}$ associated with a miner $A$ is communicated to the mining pool and subsequently assigned to an auditor in a centralized pool setting. We further discuss how analogous information can be disseminated in a decentralized mining pool, allowing multiple participants to independently reason about and audit the claimed search space $S_{i}^{A}$.
@@ 3.1 Scheme 1: Non-Empty V-Blocks
Let us assume that we associate a new block header with the PoW after it is found. Then the parent block will also be chosen post-PoW, which means that a v-block could be sealed to fit in different slots of the blockchain. This could disrupt the incentives of mining. A careful analysis is required including how this impacts mining fairness and selfish mining attacks. Without signature equivocation and penalization of a stake, generating valid v-blocks is free. To prevent attacks, we add a blockchain consensus rule that a v-block can be included only just immediately after its audited block. Without this consensus rule, the protocol would need to check each new v-block against all past v-blocks to prevent double-inclusion. The downside of the immediate inclusion rule is that the audit time is reduced to a single block interval, which can vary widely.
@@ 3.2 Scheme 2: Empty V-Blocks
When a v-block does not carry neither transaction nor block timestamp information, the v-block does not need sealing and is already unique. Although this option is possible, it has the downside that it reduces the blockchain transaction throughput. We analyze this variant in Appendix B.
@@ 3.7 Avoidance of Long Forks
Several hybrid consensus protocols that combine proof-of-work and proof-of-stake, such as 2-hop [8] and TwinsCoin [9], explicitly tolerate single-block forks by enforcing that the block following a PoS-extended block must be a standard PoW block. This structural constraint prevents the formation of long runs of stake-extended blocks and ensures that forks cannot grow unboundedly without additional PoW investment.
@@
Figure 2: A blockchain mixing blocks and v-blocks auditing previous work units (w.u.), with v-blocks immediately followed by a regular PoW block.
@@ 3.8 V-Block Attribution
Every block $i$ contains the field $\mathsf{poolAddress}_{i}$ in its block template. This is the address where v-blocks are paid. The address is extracted from the message $M$ in the v-block. The address can be controlled by the pool server, or it can be a smart contract controlled address in the mainchain (if it has this capability) or it can be a sidechain exit address that forwards the funds to a smart contract on a sidechain. The pool server or pool smart contract is responsible for sharing the reward among the workers, in the same manner as a normal block reward. The server can also track who was the auditor of the audited block and pay a special prize to that worker. Alternatively, a field $\mathsf{lastBlock}_{i}$ could be added to the block template. This field would indicate the block ID of the last block created by the same pool that was included in the canonical chain. The protocol could retrieve that block coinbase transaction and distribute the reward to the same addresses and in the same proportion used in the referenced block. This enables Bitcoin decentralized mining pools that pay their workers directly from the coinbase transaction using multiple outputs.
@@
3.9 Pool Share Difficulty Management
@@
4 APoW Compared to Prior Work
We compare our scheme with the existent solutions to BWA.
4.1 Oblivious Shares vs Our Approach
4.2 “Pop quiz” (and its variants) vs Our Approach
In pop-quiz schemes, the hashing spent verifying other miners is not rewarded by the blockchain, limiting their practical effectiveness as a BWA deterrent.
5 Security
We identify three potential attacks related to the fairness of mining and v-mining, and we show how our scheme resist such attacks.
@@ 5 Security
Memory-CPU trade-off provides little help with regard to RAM costs. Assuming that all hash operations are stored sequentially, each cached hash operation $j$ requires the storage of $X_{j}$, which is 256-bits in length. It is neither necessary nor optimal to store $X_{j}$ in full. The digests $X_{j}$ could be partially stored and recomputed in full after a positive match with the partial target pattern, to verify the full match. For example, if we store 16-bits per digest, we can retrieve 8 hash results per RAM memory access (assuming a 128-bit wide data bus). This results in an 8X speedup for RAM access and still leaves plenty of time for a CPU or a non-pipelined hashing core to recompute promising partial matches. Since the nonce candidates to check will not be sequential, they need to be transmitted to the hashing component, thus the communication can become the system’s bottleneck. While this optimization could make storing and retrieving results as fast as hashing, the v-mining machine would still cost over 1T dollars.
The attacker can also store just 1 bit per hash, packing bits in a 128-bit hint word that can be retrieved in a single RAM access. The attack requires the building of a special ASIC architecture that reads hint words, scans its bits while it increments the nonce counter at double clocking speeds. This small overclocked circuit would stall for 1 cycle when a valid bit (0 or 1) is found in the hint bit sequence, so that the corresponding nonce enters the hashing pipeline, keeping it almost full. Although this design may work, the machine would still cost 39B USD due to its large RAM requirement.
@@
Figure 3: Recursive Auditing: audit at height $j$ conducted for an auditing work unit at height $k$, which in turns audits a work unit at height $i$
@@
6 implementation
In this section, we analyze the requirements and limitations when implemeting APoW on Bitcoin.
@@ 6.3 Migrating Bitcoin PoW to APoW
While switching to a new mining function and disposing all pre-existing ASICs is unfeasible, Bitcoin could hard-fork to enable both methods of mining to co-exist. Mining pools could offer an additional incentive for miners that allow v-mining, and with enough hashrate v-mining, no miner would require to pass KYC checks or amass reputation.
@@
6.4 APoW Without Blockchain Rewards
APoW is most effective when the underlying blockchain protocol explicitly rewards the submission of verification blocks (v-blocks), such rewards are not strictly required for the construction to remain useful. Even in the absence of direct compensation for hash power devoted to auditing, a mining pool can still safely assign resources to verification tasks: shares generated during auditing are indistinguishable from ordinary shares and can therefore be accounted for under the pool’s existing payout mechanisms. As long as the expected cost of auditing remains lower than the expected loss induced by a successful block withholding attack, the pool can rationally fund audits without exposing itself to additional avenues of cheating.
6.5 Defining Auditable Work Units in Mining Pools
@@ Appendix B Unsealed V-Blocks
    Scheme 2c. The v-block header is not part of the blockchain header chain, but is included in the blockchain inside a normal transaction $T$ with a special payload (i.e., OP_RETURN data in Bitcoin, or calldata in EVM-like chains). The v-block does not contribute to the cumulative work of the chain, but the protocol mints a subsidy $S$ to reward the v-miner. The transaction $T$ has an output with value $S$, which is considered as a virtual input of the transaction, same as the coinbase.
@@
Mining v-blocks should be almost as profitable as mining normal blocks to avoid penalizing pools that use them to perform audits. However, it should not be equally profitable. If so, the mining pools could engage only in v-mining, reducing the available transaction volume on the blockchain and preventing other pools from auditing. In Bitcoin, v-mining would be discouraged as unsealed v-blocks cannot earn transaction fees.